OpenStudy (anonymous):

Help pleaseeeee!!! Directions: Write an equation to represent the given relationship among integers. Use this definition: The product of any real number and an integer is called a multiple of the real number. jim weighs more than joe but less than jack. their weights in kilograms are consecutive multiples of 7. if they each weighs 5 kg less, the sum of their weights would be 195 kg.

OpenStudy (swissgirl):

(7x-5)+(7x+7-5)+(7x+14-5)=195

OpenStudy (swissgirl):

This is the equation still working on it though

OpenStudy (swissgirl):

ok yes this is the equation. I just double checked it

OpenStudy (swissgirl):

Ok Jack weighs the most and Joe weighs the least Ok what we do know is that each one weighs 7 more than the next and they are multiples of 7 Lets start of with Joe Joe=7x Jim=7x+7 Jack= 7x+7+7=7x+14

OpenStudy (swissgirl):

Ok the next thing we know is that the sum of 5 kg less than all of their weight =195 Joe=7x-5 Jim=7x+7-5=7x+2 Jack=7x+14-5=7x+9 (7x-5)+(7x+2)+(7x+9)=195

OpenStudy (swissgirl):

21x+6=195 21x=189 x=9 Joe=7(9)=63 Jim=7(9)+7=70 Jack=7(9)+14=77

OpenStudy (anonymous):

Oh my gosh thanks soooo much!!!!!

OpenStudy (swissgirl):

you are welcome